Locate and classify the stationary points of the following:

(i) f(x,y) = 4xy + x4 - y4

(ii) f(x,y)= xy + 2/x + 4/y, x>0, y>0

1
Expert's answer
2021-06-02T11:34:26-0400

Ans:-

"(i ) \ f(x,y) = 4xy + x^4 - y^4"

Partial derivative of "\ f(x,y)" with respect to "x"


"\dfrac{\delta f}{\delta x}= 4y+4x^3"


Partial derivative of "\ f(x,y)" with respect to "y"

"\dfrac{\delta f}{\delta y}= 4x-4y^3"


For stationary points put "\dfrac{\delta f}{\delta x}=0" and "\dfrac{\delta f}{\delta y}=0"


from first equation "y=-x^3 \ \ and \ \" from second expression "x=y^3" Hence these condition are fulfilled by only one condition "(0,0)" and this point will be The Origin .


For classified the nature of stationary point


"D =\dfrac{∂^2f}{∂x^2}\times \dfrac {∂^2f}{∂y^2}−(\dfrac{∂^2f}{∂x∂y})^2"


For the stationary point "(0,0)"

"D=-16\ \ <0"

the stationary point is a saddle point.


"(ii) \ f(x,y)= xy + \dfrac{2}{x}+ \dfrac{4}{y}"

Partial derivative of "\ f(x,y)" with respect to "x"


"\dfrac{\delta f}{\delta x}=y-\dfrac{2}{x^2}"


Partial derivative of "\ f(x,y)" with respect to "y"


"\dfrac{\delta f}{\delta y}=x-\dfrac{4}{y^2}"


For stationary points put "\dfrac{\delta f}{\delta x}=0" and "\dfrac{\delta f}{\delta y}=0"


From first equation "yx^2=2" and "xy^2 = 4" Hence these condition are fulfilled by only one condition and this point will be "(1,2)"


For classified the nature of stationary point

"D =\dfrac{∂^2f}{∂x^2}\times \dfrac {∂^2f}{∂y^2}−(\dfrac{∂^2f}{∂x∂y})^2"


For the stationary point "(1,2)"


"\dfrac{∂^2f}{∂x^2}=\dfrac{4}{x^3} \ \ , \ \ \dfrac{∂^2f}{∂y^2}=\dfrac{8}{y^3} \ \ and\ \ \dfrac{∂^2f}{∂x∂y}=1"


"D=3>0 \ \ \ and \ \ \dfrac{∂^2f}{∂x^2}>0"

Then the stationary point is a local minimum.